The actuary for the pension plan of Bridgeport Inc. calculated the following net gains and losses.

Incurred during the Year

(Gain) or Loss

2020

$301,000

2021

475,500

2022

(209,400)

2023

(287,600)


Other information about the company’s pension obligation and plan assets is as follows.

As of January 1,

Projected Benefit
Obligation

Plan Assets
(market-related asset value)

2020

$3,968,000 $2,421,100

2021

4,529,000 2,190,800

2022

4,987,000 2,586,100

2023

4,202,300 3,050,400


Bridgeport Inc. has a stable labor force of 400 employees who are expected to receive benefits under the plan. The total service-years for all participating employees is 6,000. The beginning balance of accumulated OCI (G/L) is zero on January 1, 2020. The market-related value and the fair value of plan assets are the same for the 4-year period. Use the average remaining service life per employee as the basis for amortization.

Compute the minimum amount of accumulated OCI (G/L) amortized as a component of net periodic pension expense for each of the years 2020, 2021, 2022, and 2023. Apply the “corridor” approach in determining the amount to be amortized each year. (Round answers to 0 decimal places, e.g. 2,500.)
Answer #1

Year

Minimum Amortization of (Gain) Loss

2020

0

2021

0

2022

$18460

2023

$8501

Amortization of excess of the cumulative net gain or loss over the corridor amount = excess of the cumulative net gain or loss over the corridor amount/ average remaining service period of employees

Average remaining service life per employee= Expected future years of service/ Number of employees = 6000/400 = 15

Year

Projected Benefit Obligation

Plan Assets

Corridor

Accumulated OCI (G/L)

Minimum Amortization of (Gain) Loss

2020

3968000

2421100

396800

0

0

2021

4529000

2190800

452900

301000

0

2022

4987000

2586100

498700

776500

18520

2023

4202300

3050400

420230

548580

8557

corridor = 10% of the greater of projected benefit obligation or plan assets

Minimum Amortization of (Gain) Loss (2022)= (776500-498700)/15 =18460

Accumulated OCI (G/L) (2023) = 776500-18520-209400 = 547740

Minimum Amortization of (Gain) Loss (2023) = (548580-420230)/15 = 8557